OpenStudy (anonymous):

write the equation in point-slope form. the line perpendicular to 2x-4y=8 passing through the point (1,-2)

OpenStudy (anonymous):

first you need the slope. the slope of \[2x-4y=8\] is \[\frac{1}{2}\] so slope of perpendicular line will be -2

OpenStudy (anonymous):

then use 'point slope" formula \[y-y_1=m(x-x_1)\] with \[m=-2, x_1=-,y_1=-2\] to get \[y+2=-2(x-1)\]

OpenStudy (anonymous):

multiply and get \[y+2=-2x+2\] \[y=-2x\]